Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum.

Answer by checkley71(8403)

3B+G=175 OR G=175-3B
2B+4G=200 NOW SUBSTITUTE (175-3B) FOR G IN THE SECOND EQUATION WE GET
2B+4(175-3B)=200
2B+700-12B=200
-10B=-500
B=-500/-10
B=50 SOLUTION ($.50 FOR THE BARS)
G=175-3*50
G=175-150
G=25 SOLUTION ($.25 FOR THE GUM)
PROOF
2*50+4*25=200
100+100=200
200=200